Prepare a cost exhibit showing cumulative cost by time period for the project shown below.

Activity

Immediate Predecessors

Duration

Activity Budget Cost

1.0

4

1,000

2.1

1.0

3

6,000

2.2

1.0

4

5,000

2.3

1.0

2

1,500

3.0

2.1

4

3,200

4.0

2.2 2.3

3

4,800

5.0

3.0   4.0

4

800

6.1

5.0

3

7,200

6.2

5.0

2

1400

6.3

5.0

1

1,200

7.0

6.1 6.2 6.3

2

2,100

  1. What is the budgeted cumulative cost at the end of period 5 based on Early Start?   
  2. What is the budgeted cumulative cost at the end of period 11 based on Early Start?    
  3. During project execution, activity 5.0 took 1 extra period with an extra cost of 300 and activity 6.3 took an extra period with an extra cost of 400. What is the schedule variance when activity 5.0 finishes? Include sign in your answer.    
  4. Using the information from c above, what is the cost variance when activity 5.0 finishes? Include sign in your answer.   
  5. At the conclusion of activity 5.0, what is the value of the Cost Performance Index to 4 digits?   
  6. Assuming no other new costs beyond the amounts in part c what is the final cost and duration of the project? What is the cost variance and schedule variance of the project at this point?   
Answer #1

Let's create the table and the gantt chart first

a)

At the end of period 5 the budgeted cost should cover

100% completion of activity 1.o

33.3% completion of activity 2.1

25% completion of activity 2.2

50% completion of activity 2.3

Thus the budgeted cost is 1000 + (1/3)*6000 + (1/4)*5000 + (1/2)*1500 = 5000

b)

At the end of period 11 the following is budgeted to be completed. Activities 1.0, 2.1, 2.2, 2.3, 3.0, and 4.0. Thus the total cost is 21500

c)

Activity 5 is supposed to finish at the end of period 15. This means the budgeted cost of work scheduled is 22300 at this point. However, if it takes additional 1 period then it will complete at end of period 16.

Budgeted cost of work scheduled (BCWS) at the end of period 16 is

22300 + (1/3)*7200 + (1/2)*1400 + 1200 = 26600

However, the Budgeted cost of work performed (BCWP) is only till activity 5. That means it is 22300.

The formula for schedule variance is

SV = BCWP – BCWS = 22300 – 26600 = -4300

d)

Actual cost of work performed (ACWP) incurred is 22300+300 = 22600

The formula for cost variances

CV = BCWP – ACWP = 22300 – 22600 = -300